# Quantifying the Dunkelflaute: An analysis of variable renewable energy droughts in Europe

### v1

@@ 2.2 Maximum drought duration strongly depends on threshold
Figure 2 shows the longest droughts obtained from the data sorted by year. A general finding is that maximum drought duration strongly depends on the underlying drought threshold. For very low thresholds around $\tau=0.1$, only very short droughts can be found. For high thresholds approaching $\tau=1$, the drought duration grows strongly and reaches a plateau of 365 days, i.e., a full year, in most weather years. This is because the search algorithm identifies very long events that reflect below-average yearly renewable availability at very high thresholds. Droughts that are relevant for energy system operations are likely to be found between these two extremes. Using very high thresholds close to $\tau=1$ thus appears not to be meaningful as these are likely to identify below-average weather years rather than drought events within a given year.
For medium thresholds, we find that both a technology portfolio and geographical balancing strongly decrease the maximum drought duration. For example, for a threshold of $\tau=0.75$, the longest PV drought in Germany lasts 291 days, while the maximum on- and offshore wind droughts are 303 and 209 days, respectively. In the case of a combined VRE portfolio, the respective maximum drought in Germany decreases to 106 days (portfolio effect). If, on top, perfect interconnection in Europe is considered, the duration of the longest event for the same threshold shrinks to only 55 days (balancing effect).
Especially for medium thresholds, the maximum yearly drought duration strongly depends on the weather year. For example, the longest VRE portfolio drought in Germany obtained for a threshold of $0.75$ occurred in 1996 (106 days), while the shortest yearly maximum drought duration was in 2018 (22 days). Considering perfect interconnection across all countries and the same threshold, the longest and shortest portfolio droughts occurred in 1997 and 1999, respectively (55 and 8 days). The most extreme droughts vary substantially across countries, both with respect to their duration and the year of occurrence (Figures SI.5 and SI.6). Importantly, the ranking of years varies with the drought threshold (Figure 3). For instance, the longest VRE portfolio drought in Germany for a threshold of 0.4 occurred in 2007 but for the thresholds 0.6 and 0.8 the most extreme droughts were in 1997 and 2003. The corresponding longest drought events in Spain occurred in 1991, 2017, and 2017. The spreads between the longest drought events in each year further increase with higher thresholds, before they decrease again approaching the threshold $\tau=1$ (Figure SI.7). From these findings follows that any claim on a specific year being particularly relevant for renewable drought analyses should always be qualified with the respective threshold.
@@
Figure 3: Most extreme duration of single drought events per year and threshold. The year with the most extreme event duration varies across thresholds.
@@
The maximum drought duration varies substantially across seasons and technologies. Solar PV droughts primarily occur in fall and winter, with their median hour in November, December, or January (Figure 4). The maximum European PV drought under unconstrained geographical balancing lasts 248 days with a median hour in December for a threshold of $\tau=0.75$. The longest PV droughts with median hours in spring or summer months barely last longer than a few days, except for those identified by very high thresholds. In contrast, the longest European on- and offshore wind droughts mainly occur in spring and summer, with medians in July and maximum duration of 218 or 204 days, respectively. Maximum wind drought duration in winter months are much lower. In single countries, summer and winter wind droughts can be significantly longer (balancing effect). The seasonal complementarity of wind and solar further mitigates combined technology portfolio droughts to a substantial extent (portfolio effect). The longest-lasting droughts occur during winter months, which is a result of low seasonal PV availability in combination with rare but severe concurrent wind droughts. The median hours of the longest technology portfolio droughts in Germany (106 days) or Spain (137 days) occur in November, or, under the assumption of unconstrained balancing, in January (55 days).
@@
Figure 4: Exemplary most extreme duration of single drought events across all months and thresholds. The contour lines represent the threshold-month-specific maximum duration. Note that the monthly assignment illustrates the median hour of respective droughts, while the duration of each event is plotted on the vertical axis. Events lasting longer than one month start (end) in previous (subsequent) months. The events with the highest and lowest duration across all months are marked for a threshold $\tau=0.75$.
@@
Comparing weighted averages over all thresholds and countries, the maximum drought duration of a VRE portfolio decreases by 64%, 52%, or 47% compared to solar PV, onshore wind, or offshore wind alone (Table 1). The portfolio effect is more pronounced for PV than for wind power, as onshore and offshore wind power complement both diurnal and seasonal shortages of PV. In contrast, wind power does not incur diurnal shortages. Assuming a European copperplate (CP), the portfolio effect (-80%) for solar PV is even stronger than for any isolated country. This is because in a larger geographical area, there is more complementary wind power available to compensate for solar PV shortages [roth_geographical_2023]. In turn, the portfolio effect is less pronounced for onshore wind power in the European interconnection (-45%) as PV availability is more homogeneously distributed. For offshore wind power, the portfolio effect is higher (-70%), as only a subset of countries features this technology, such that solar and wind onshore resources of all countries can compensate for extreme offshore droughts in a subset of countries.
@@
Considering unconstrained geographical balancing across all countries, the longest PV, onshore wind, or offshore wind droughts shorten by 1%, 46%, or 34%, using weighted averages over all countries. The very small average balancing effect for PV, however, conceals that there is a substantial north-south divide between countries with higher and lower solar irradiation. For South European countries with abundant solar availability such as Spain, Portugal, Italy, or Greece, geographical balancing even increases solar PV drought duration. In turn, geographical balancing alleviates maximum PV drought durations in North European countries such as Germany or Scandinavia. In contrast, the balancing effect on wind drought duration is more similarly distributed across countries, as maximum wind droughts rarely occur simultaneously throughout Europe. The balancing effect is even more pronounced for VRE technology portfolios, decreasing by 65% on average. That is, geographical balancing not only helps to mitigate droughts for single renewable technologies, but even more so for a technology portfolio, due to complementary technology portfolios with different capacity mixes across countries.
@@
2.3 Multi-threshold perspective reveals temporal dynamics of drought patterns
The results discussed so far reveal a strong sensitivity of drought characteristics to the underlying threshold used for drought identification. Focusing on events lasting at least two days, Figure 5 exemplarily illustrates drought patterns for a wide range of thresholds for the years 1996 and 1997. In general, single-threshold analyses using low thresholds can detect extreme short-duration droughts, which may occur in isolation or adjacent to high-availability periods (white spaces in the Figure). Higher thresholds increasingly identify longer-lasting events. These potentially include consecutive shorter periods of extreme renewable shortage or surplus, which are smoothed by the averaging mechanics of the identification algorithm (see Section 4.1). The multi-threshold perspective illustrated in Figure 5 demonstrates that longer-lasting events identified by higher thresholds may encompass multiple severe droughts detected at lower thresholds.
Figure 5: Identified drought patterns in 1996 and 1997 for all investigated thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one region. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Figure SI.8 focuses on longer-lasting events by illustrating respective patterns for droughts lasting at least one week.
Solar PV events identified by moderate or high thresholds can last several months, indicating low solar availability during winter in Europe (Figure 5, panel b). Lower thresholds identify very severe PV droughts driven by low seasonal availability in combination with factors such as persistent cloud or snow coverage, such as in winter 1996. In summer, only rare and isolated PV droughts of lower severity occur. These seasonal PV characteristics can generally be found across all years and are often more pronounced in Northern Europe, e.g., observable in winter 2011/12 (compare Germany or United Kingdom to Spain in Figure SI.9). In contrast, severe long-lasting onshore and offshore wind droughts can occur throughout the year but tend to be more frequent in summer (Figure 5, panels c and d), which is in line with general wind seasonality [kaspar_climatological_2019]. The multi-threshold perspective further reveals that severe wind droughts identified by lower thresholds can be sequential and may occur within contiguous below-average wind periods detected at higher thresholds. The latter may last up to several months. While such long-lasting events usually occur in summer (compare Figures SI.9-SI.11), they may also take place in winter in some countries (e.g., in 1996/97, Figure 5). On- and offshore droughts are generally correlated, but may differ in severity (Figure 5, panels c and d).
The multi-threshold illustration also shows that wind and solar power droughts are often complementary. Typically, high solar availability offsets wind droughts in summer, while wind power compensates for the low solar PV availability in winter. This results not only in briefer but also less severe droughts, i.e., higher thresholds apply in the same periods compared to single technologies (portfolio effect, compare panel a of Figure 5 with panels b-d). However, if severe and persistent wind and PV droughts coincide, pronounced long-lasting portfolio droughts can occur, as seen for instance in winter 1996/97. Geographical balancing can mitigate the severity and duration of such compound events as solar seasonality is lower in Southern Europe and wind droughts do not occur simultaneously across countries (balancing effect). Severe and long-lasting portfolio droughts may arise in single countries and also across Europe at the beginning of a calendar year, (e.g., observable in 2012, Figure SI.9), at the end of a year (e.g., 1982, Figure SI.11), or span across the turn of years (e.g., in winter 1996/97, Figure 5). In some years, winter portfolio droughts are much less pronounced (e.g., 2013/14, Figure SI.10). Countries relying heavily on wind power, such as the United Kingdom, may also experience very severe events in summer.
@@ 2.3 Multi-threshold perspective reveals temporal dynamics of drought patterns
2.4 Drought mass: a new metric to identify drought events relevant for long-duration electricity storage
We introduce a novel “drought mass” metric to identify the most extreme portfolio drought events and show that it identifies events that are relevant for long-duration electricity storage in renewable European energy systems. The drought mass rates both the duration and severity of contiguous events by accumulating and ranking scores for all droughts identified at thresholds $\tau\leq 0.75$. Section 4.2 further elaborates on the drought mass mechanics.
@@ 2.4 Drought mass: a new metric to identify drought events relevant for long-duration electricity storage
Figure 6 shows extreme portfolio drought patterns identified by the drought mass metric for the years 1996/97. The most extreme events, i.e., those with the highest drought mass, comprise sequences of shorter, but more severe droughts within contiguous well-below-average periods that may last up to several months. Events with the highest drought mass scores may occur in winter (teal boxes), potentially spanning across the turn of a calendar year, as in the case of the European copperplate in 1996/97. Severe drought mass events may also occur in summer (gray boxes) in countries that heavily rely on wind power such as the United Kingdom or Poland.
@@
Figure 6: Identified most extreme drought events in 1996/97 occurring in winter (teal boxes). For countries in which the most extreme drought events occur in summer, these are additionally shown (gray boxes). For each region, portfolio drought occurrences lasting longer than one day for color-coded thresholds (upper panel) as well as normalized exogenous demand profiles and optimized storage levels across three modeled interconnection scenarios (lower panel) are displayed.
Peak electricity demand often depends on ambient temperature and occurs in winter in most European countries, except for the Mediterranean area. Accordingly, winter drought events that coincide with high electricity demand are particularly relevant. Using a stylized electricity sector model (Section 4.3), we find that these compound events are a major driver for the use of long-duration electricity storage in all three interconnection states. Figure 6 shows the optimal operation of long-duration electricity storage for different degrees of interconnection between countries as well as electricity demand patterns for the years 1996/97. The most severe identified droughts largely coincide with the periods of long-duration storage discharge, i.e., decreasing storage levels. This holds true for both the European copperplate scenario and several isolated countries in the island scenario. When considering policy-relevant interconnection across all countries (“TYNDP”), the storage level patterns do not structurally change but only show a level shift compared to the island scenarios. That is, drought patterns identified for isolated countries are a reasonable approximation of drought patterns in settings with more policy-relevant interconnection with regard to long-duration storage operation.
In some countries, the most severe drought mass event does not coincide with the major discharging period of long-duration storage. For example, in Sweden, Poland, or the United Kingdom, the metric finds that the largest drought events occur in summer (gray boxes in the lowest three panels of Figure 6). Long-duration storage usage, however, still coincides with the most extreme winter droughts, which are comparatively less severe. This is because electricity demand is lower during the more severe summer drought events identified in these countries. Consequently, summer droughts do not affect the system’s ability to meet demand as much. In complementary model runs where we assume a flat electricity demand profile, i.e., abstracting from seasonal and diurnal variability, the most severe event measured by the drought mass metric perfectly coincides with storage-defining periods in all countries, including Sweden, Poland, and the United Kingdom (Figure SI.12).
@@ 2.5 Geographical balancing inside the European renewable energy super drought
Figure 8 shows four snapshots of selected hours within the European portfolio super drought, as identified by the drought mass metric at the turn of the years 1996/97. Each panel shows the drought severity present in the displayed hour for isolated countries (island scenario, left map) and for the European copperplate scenario (right map), illustrated by the lowest applying threshold. While the super drought affects multiple European countries simultaneously, the severity varies across countries and over time. Comparing the island scenario and the European copperplate shows that renewable energy shortages in the latter are always less severe than in the most affected countries at the same hour. This is because the assumed European interconnection mitigates shortages in one country by leveraging higher renewable availability in others. That is, geographical balancing is feasible even within the most severe pan-European drought event in the data. The graphs further illustrate that the super drought comprises sequences of shorter, but more severe droughts, both in individual countries (different colors on maps on the left-hand side) and on the European level (different colors between maps on the right-hand side). Further, even within the European super drought, there are hours with relatively high renewable availability (Figure 8d). The progression of sequentially occurring events inside the European super drought is even more visible in a complementary animated graph which is available as supplementary information [kittel_high_2024]. The animation, covering December 1996 and January 1997, shows that very severe droughts that cover most of Europe (e.g., visible in Figure 8a) are relatively brief and occur rarely, even inside the super drought event. That is, some renewable generation potential is always available somewhere in Europe. Even during the most extreme droughts, this can balance low renewable availability across technologies and regions to some extent.
(a) Severe droughts covering almost all of Europe.
(b) Severe droughts mainly in Central Europe.
(c) Severe droughts covering almost all of Continental Europe.
(d) Moderate droughts in most of Europe.

### v2

@@ 2.3 Maximum drought duration strongly depends on threshold and differs across years and seasons
Figure 3 shows the longest droughts obtained from the data across all investigated years for different relative thresholds and technologies. A general finding is that the maximum drought duration strongly depends on the threshold. For very low thresholds around $\tau=0.1$, the maximum duration of all droughts in the data is very short. For higher thresholds approaching $\tau=1$, the maximum drought duration grows strongly and eventually reaches a plateau of 365 days, i.e., a full year. This is because at very high thresholds, the search algorithm identifies years with below-average renewable availability as drought events. Using very high thresholds close to $\tau=1$ thus appears not to be meaningful for identifying droughts relevant for energy system operations within a given year.
@@
Figure 3: Most extreme duration of single drought events across all investigated years for given relative thresholds $\tau\in[0.1,1]$. The year with the most extreme event duration varies across thresholds. The marked numbers indicate the longest drought durations found for a relative threshold of $\tau=0.75$.
For medium thresholds, we find that both a technology portfolio and geographical balancing strongly decrease the maximum drought duration. For example, for a relative threshold of $\tau=0.75$, the longest PV drought in Germany lasts 291 days, while the maximum on- and offshore wind droughts are 303 and 209 days, respectively. In the case of a combined VRE portfolio, the respective maximum drought in Germany decreases to 106 days due to the technology portfolio effect. If, on top, perfect interconnection in Europe is considered, the duration of the longest portfolio event for the same threshold further declines to 55 days due to the geographical balancing effect.
Especially for medium thresholds, the maximum yearly drought duration strongly depends on the weather year (Figure SI.9). For example, the longest VRE portfolio drought in Germany obtained for a relative threshold of $\tau=0.75$ occurred in 1996 (106 days), while the shortest yearly maximum drought duration was in 2018 (22 days). Considering perfect interconnection across all countries and the same threshold, the longest and shortest portfolio droughts occurred in 1997 and 1999, respectively (55 and 8 days).
The most extreme droughts vary substantially across countries, both with respect to their duration and the year of occurrence (Figures SI.10 and SI.11). Importantly, the ranking of years varies with the drought threshold (Figure 3). For instance, the longest VRE portfolio drought in Germany for a threshold of $\tau=0.4$ occurred in 2007 but for the thresholds $\tau=0.6$ and $\tau=0.8$ the most extreme droughts were in 1997 and 2003. The corresponding longest drought events in Spain (or in Europe, assuming perfect interconnection) occurred in 1991 (2005), 2017 (1996), and 2017 (2002). The spreads between the longest drought events in each year further increase with higher thresholds, before they decrease again approaching the threshold $\tau=1$ (Figure SI.12). From these findings follows that any claim on a specific year being particularly relevant for renewable drought analyses should always be qualified with the respective threshold.
@@
The maximum drought duration further varies substantially across seasons and technologies. Solar PV droughts primarily occur in fall and winter, with their median hour in November, December, or January (Figure 4). The maximum European PV drought under unconstrained geographical balancing lasts 248 days with a median hour in December for a threshold of $\tau=0.75$. The longest PV droughts with median hours in spring or summer months barely last longer than a few days, except for those identified by very high thresholds close or equal to $\tau=1$. In contrast, the longest European on- and offshore wind droughts mainly occur in spring and summer, with medians in July and maximum duration of 218 or 204 days, respectively. Maximum wind drought duration in winter months are much lower. In single countries, summer and winter wind droughts can be significantly longer (balancing effect). The seasonal complementarity of wind and solar further mitigates combined technology portfolio droughts to a substantial extent (portfolio effect). The longest-lasting droughts occur during winter months, which is a result of low seasonal PV availability in combination with rare but severe concurrent wind droughts. The median hours of the longest renewable technology portfolio droughts in Germany (106 days) or Spain (137 days) occur in November, or, under the assumption of unconstrained balancing, in January (55 days).
@@
Figure 4: Example of most extreme duration of single drought events across all months in the data and all investigated thresholds $\tau\in[0.1,1]$. The contour lines represent the threshold-month-specific maximum duration. Note that the monthly assignment illustrates the median hour of respective droughts, while the duration of each event is plotted on the vertical axis. Events lasting longer than one month start (end) in previous (subsequent) months. The events with the highest and lowest duration across all months are marked for a threshold $\tau=0.75$.
2.4 Portfolio and balancing effects mitigate maximum drought durations
@@
Comparing weighted averages over all thresholds and countries, the maximum drought duration of a VRE portfolio decreases by 64%, 52%, or 47% compared to solar PV, onshore wind, or offshore wind alone (Table 1). This portfolio effect is more pronounced for PV than for wind power, as onshore and offshore wind power complement both diurnal and seasonal shortages of PV. In contrast, wind power does not incur diurnal shortages. Assuming a European copperplate (CP), the portfolio effect (-80%) for solar PV is even stronger than for any isolated country. This is because in a larger geographical area, there is more complementary wind power available to compensate for solar PV shortages [roth_geographical_2023]. In turn, the portfolio effect is less pronounced for onshore wind power in the European interconnection (-45%) as PV availability is more homogeneously distributed. For offshore wind power, the portfolio effect is higher (-70%), as only a subset of countries features this technology, such that solar and wind onshore resources of all countries can compensate for extreme offshore droughts in a subset of countries.
@@
Considering unconstrained geographical balancing across all countries, the longest PV, onshore wind, or offshore wind droughts decrease by 1%, 46%, or 34%, using weighted averages over all countries. The very small average balancing effect for PV, however, conceals that there is a substantial north-south divide between countries with higher and lower solar irradiation. For South European countries with abundant solar availability such as Spain, Portugal, Italy, or Greece, geographical balancing even increases solar PV drought duration. In turn, geographical balancing alleviates maximum PV drought durations in North European countries such as Germany or Scandinavia. In contrast, the balancing effect on wind drought duration is more similarly distributed across countries, as maximum wind droughts rarely occur simultaneously throughout Europe. The balancing effect is even more pronounced for VRE technology portfolios, decreasing by 65% on average. That is, geographical balancing not only helps to mitigate droughts for single renewable technologies, but even more so for a technology portfolio, due to complementary technology portfolios with different capacity mixes across countries.
@@
2.5 Drought mass: a multi-threshold metric to identify drought events relevant for long-duration electricity storage
Previous studies that use threshold-based identification methods typically use only a limited number of thresholds (e.g., [leahy_persistence_2013, cannon_using_2015, patlakas_low_2017, kaspar_climatological_2019, ohlendorf_frequency_2020, kies_critical_2021, potisomporn_extreme_2024, potisomporn_extreme_2024, abdelaziz_assessing_2024, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022]). However, the findings presented above indicate that drought characteristics strongly depend on the chosen threshold. To account for this, we introduce a multi-threshold “drought mass” metric to identify the most extreme portfolio drought events and show that it identifies events that are relevant for long-duration electricity storage in renewable European energy systems. In contrast to previous single-threshold approaches, the drought mass quantifies extreme events by integrating both drought duration and severity across a wide range of thresholds, thus properly accounting for the complex and diverse renewable drought patterns discussed in Section 2.1. It accumulates the number of drought hours identified at thresholds $\tau\leq 0.75$ that occur within contiguous events identified at $\tau=0.75$, i.e., excluding droughts found by very high thresholds ($\tau>0.75$). The event with the highest score within a each pair of consecutive years in the data (i.e., 1982-1983, 1983-1984, 1984-1985, etc.) is identified as the most extreme drought event within this time frame. Section 4.2 further elaborates on the drought mass mechanics.
@@
Figure 5 shows extreme portfolio drought patterns identified by the drought mass metric for the years 1996/97 and selected countries. The most extreme events, i.e., those with the highest drought mass, comprise sequences of shorter, but more severe droughts within contiguous well-below-average periods that may last up to several months. Events with the highest drought mass scores may occur in winter (teal boxes), potentially spanning across the turn of a calendar year, as in the case of the European copperplate in 1996/97. Severe drought mass events may also occur in summer (gray boxes) in countries that heavily rely on wind power such as the United Kingdom or Poland.
@@
Figure 5: Most extreme drought events in 1996/97 of selected countries identified by the drought mass metric occurring in winter (teal boxes). For countries in which the most extreme drought events occur in summer, these are additionally shown (gray boxes). For each region, portfolio drought occurrences lasting longer than one day for color-coded thresholds (upper panel) as well as normalized exogenous demand profiles and optimized storage levels across three modeled interconnection scenarios (lower panel) are displayed.
Peak electricity demand often depends on ambient temperature and occurs in winter in most European countries, except for the Mediterranean area. Accordingly, winter drought events that coincide with high electricity demand are particularly relevant. Using a stylized electricity sector model (Section 4.3), we find that these compound events are a major driver for the use of long-duration electricity storage in all three interconnection states. Figure 5 shows the optimal operation of long-duration electricity storage for different degrees of interconnection between countries as well as electricity demand patterns for the years 1996/97. The most severe identified droughts largely coincide with the periods of long-duration storage discharge, i.e., decreasing storage levels. This holds true for both the European copperplate scenario and several isolated countries in the island scenario. When considering policy-relevant interconnection across all countries (“TYNDP”), the storage level patterns do not structurally change but only show a level shift compared to the island scenarios. That is, drought patterns identified for isolated countries are a reasonable approximation of drought patterns in settings with more policy-relevant, limited interconnection with regard to long-duration storage operation.
In some countries, the most severe drought mass event does not coincide with the major discharging period of long-duration storage. For example, in Sweden, Poland, or the United Kingdom, the metric finds that the largest drought events occur in summer (gray boxes in the lowest three panels of Figure 5). Long-duration storage usage, however, still coincides with the most extreme winter droughts, which are comparatively less severe. This is because electricity demand is lower during the more severe summer drought events identified in these countries. Consequently, summer droughts do not affect the system’s ability to meet demand as much. In complementary model runs where we assume a flat electricity demand profile, i.e., abstracting from seasonal and diurnal variability, the most severe event measured by the drought mass metric perfectly coincides with storage-defining periods in all countries, including Sweden, Poland, and the United Kingdom (Figure SI.13).
@@ 2.6 Geographical balancing inside the European renewable energy super drought
Figure 7 shows four snapshots of selected hours within the European portfolio super drought, as identified by the drought mass metric at the turn of the years 1996/97. Each panel shows the drought severity present in the displayed hour for isolated countries (island scenario, left map) and for the European copperplate scenario (right map), illustrated by the lowest applying threshold. While the super drought affects multiple European countries simultaneously, the severity varies across countries and over time. Comparing the island scenario and the European copperplate shows that renewable energy shortages in the latter are always less severe than in the most affected countries at the same hour. This is because the assumed European interconnection mitigates shortages in one country by leveraging higher renewable availability in others. That is, geographical balancing is feasible even within the most severe pan-European drought event in the data. The graphs further illustrate that the super drought comprises sequences of shorter, but more severe droughts, both in individual countries (different colors on maps on the left-hand side) and on the European level (different colors between maps on the right-hand side). Further, even within the European super drought, there are hours with relatively high renewable availability (Figure 7d). The progression of sequentially occurring events inside the European super drought is even more visible in a complementary animated graph which is available as supplementary information [kittel_high_2025]. The animation, covering December 1996 and January 1997, shows that very severe droughts that cover most of Europe (e.g., visible in Figure 7a) are relatively brief and occur rarely, even inside the super drought event. That is, some renewable generation potential is always available somewhere in Europe. Even during the most extreme droughts, this can balance low renewable availability across technologies and regions to some extent.
(a) Severe droughts covering almost all of Europe on 7 December 1996 at 8pm.
(b) Severe droughts mainly in Central Europe on 19 December 1996 at 8am.
(c) Severe droughts covering almost all of Continental Europe on 10 January 1997 at midnight.
(d) Moderate droughts in most of Europe on 13 January 1997 at noon.